# Overcoming the Limitations of Layer Synchronization in Spiking Neural Networks

### v1

@@ A.1.1 Incoming current
where $s_{j}^{(l)}[t]$ is $1$ if neuron $j$ in layer $l$ has emitted a spike at time $t$, otherwise 00.
@@ A.1.2 Membrane potential decay
where $\tau_{m}$ is the membrane time constant and $\Delta$t the elapsed time.
@@ A.4 Class prediction
where $c_{i}$ is the spike count for class $i$, $T$ are all the timesteps for which a forward pass occurred, and $s[i,t]$ is the output of the neuron representing class $i$ in the output layer at the end of the forward pass at time $t$. These values are subsequently used as logits within a softmax function:
@@ A.5 Parameters for the simulator
Barrier-based / layer Loihi [10], SENECA [62], POETS [50] # neu/layer # neu/layer [True]
Timer-based / core SENECA [62], SpiNNaker [39], TrueNorth [1], POETS [50] # neu/core # neu/core [True]
Asynchronous Speck [6], $\upmu$Brian [53] # neu/layer 1 [False]
@@ A.6.1 Datasets and network architectures
The Neuromorphic MNIST (N-MNIST) dataset captures the MNIST digits using a Dynamic Vision Sensor (DVS) camera. It presents minimal temporal structure [20]. It consists of 60000 training samples and 10000 test samples. Each sample spans approximately 300 ms, divided into three 100 ms camera sweeps over the same digit. Only the initial 100 ms segment of each sample is used in this study.
@@
The Spiking Heidelberg Digits (SHD) dataset [8] is composed of auditory recordings with significant temporal structure. It consists of 8156 training samples and 2264 test samples. Each sample includes recordings of 20 spoken digits transformed into spike sequences using a cochlear model, capturing the rich dynamics of auditory processing. The 700 cochlear model output channels are downsampled to 350 channels.
@@
The DVS gesture dataset [2] focuses on different hand and arm gestures recorded by a DVS camera. Like the SHD dataset, it has significant temporal structure. It focuses on 11 different hand and arm gestures recorded by a DVS camera. It consists of 1176 training samples and 288 test samples. The $128\times 128$ input frame is downsampled to a $32\times 32$ frame.
For all three datasets, input events are assigned a timestamp and an index. In the case of N-MNIST, the index corresponds to a position within a $34\times 34$ pixel frame, with each pixel having a binary polarity value (either 1 or 0), leading to a total of $34\times 34\times 2=2312$ distinct input indices. For SHD, the index denotes one of the 350 output channels. For DVS gesture, the $128\times 128$ input frame is downsampled to a $32\times 32$ frame. Like N-MNIST, each pixel has a binary polarity value, so in total this gives $32\times 32\times 2=2048$ input indices.
@@
Table 5: Network architecture and hyperparameters. The architecture is given as [input features] - [neurons in hidden layers $\times$ number of hidden layers] - [neurons in output layer].
N-MNIST SHD DVS gesture
Architecture 2312-[64$\times$3]-10 350-[128$\times$3]-20 2048-[128$\times$3]-11
Timestep size 10 ms 10 ms 20 ms
Batch size 256 32 32
Epochs 50 100 70
Learning rate 5e-4 7e-4 1e-4
Membrane potential threshold $U_{\text{thr}}$ 0.3 0.3 0.3
Weight decay constant $\lambda_{W}$ 1e-5 1e-4 1e-5
Membrane time constant $\tau_{m}$ 1 ms 100 ms 100 ms
Surrogate steepness constant $\alpha$ 2 10 10
Input spike dropout 0.25 0.2 0.2
Forward group size $F$ 8 8 8
Refractory dropout 0.8 0.8 0.8
Momentum noise $\lambda_{\text{MS}}$ 1e-6 0.1 0.1

### v2

@@ A.1.1 Incoming current
where $s_{j}^{(l)}[t]$ is $1$ if neuron $j$ in layer $l$ has emitted a spike at time $t$, otherwise 0.
@@ A.1.2 Membrane potential decay
[…]
@@ A.2 Event-driven state update rule for LIF neuron
Note that the evaluation of the firing threshold takes place independently for every incoming synaptic current (instead of synchronizing and accumulating all currents first, and then evaluating the threshold). This allow to consider and preserve the temporal dynamics among the input currents (different arrival times) in the temporal derivative of the membrane potential of each neuron and propagate them accordingly.
@@ A.4 Class prediction
where $c_{i}$ is the spike count for class $i$, $T$ are all the timesteps for which a forward pass occurred, and $s[i,t]$ is the output of the neuron representing class $i$ in the output layer at the end of the forward pass at time $t$. For CIFAR-10, instead the value of $c_{i}$ is equal to the membrane potential of the output neuron for class $i$ at the end of having processed all the spikes. These values are subsequently used as logits within a softmax function:
@@ A.5 Parameters for the simulator
Barrier-based / layer Loihi Davies et al. (2018), SENECA Yousefzadeh et al. (2022), POETS Shahsavari et al. (2021) # neu/layer # neu/layer [True]
Timer-based / core SENECA Yousefzadeh et al. (2022), SpiNNaker Painkras et al. (2013), TrueNorth Akopyan et al. (2015), POETS Shahsavari et al. (2021) # neu/core # neu/core [True]
[…]
A.5.1 Neuron execution scheduling policy & asynchronous processing dynamics
A scheduling policy in the simulation environment aims to capture the biases introduced by the operating principle of an AI accelerator (class) on data dynamics, and provide a means to reflect them in the model training process (see Unlayered Backpropagation). We experimented with two scheduling policies that capture two different aspect of order dynamics, aiming to highlight their effects and importance but they are not the only possible. It is a topic of open exploration for the future, and co-design for dataflow AI accelerators.
  • Random Scheduling, reflects the fact that the spike propagation and neuron integration process is or can be inherently noisy (either epistemically or aleatorically), thus affecting the temporal/rank order of spikes. It can be seen as annealing noise or dropout noise, which beyond a certain level breaks the system down but in small enough quantities (relevant to the vector pipelining of an accelerator – $F$-group size, arbitration between cores, network-on-chip routing, etc), makes the system more robust. Note that it is applied across the entire network, not within each layer.
  • Momentum Scheduling, puts emphasis on the relative order dynamics in neuron evaluation as reflected in their membranes (i.e. which neuron is likely to fire first next – relevant to TTFS and rank-order codes for the importance of the first spike(s) even though we work with rate codes). Note again this takes place across all layers, and remains unbiased by artificial layer-synchronization barriers.
@@ A.6.1 Datasets and network architectures
The Neuromorphic MNIST (N-MNIST) dataset captures the MNIST digits using a Dynamic Vision Sensor (DVS) camera. It presents minimal temporal structure Iyer et al. (2021). It consists of 60000 training samples and 10000 test samples. Each sample spans approximately 300 ms, divided into three 100 ms camera sweeps over the same digit. Only the initial 100 ms segment of each sample is used in this study.
@@
The Spiking Heidelberg Digits (SHD) dataset Cramer et al. (2020) is composed of auditory recordings with significant temporal structure. It consists of 8156 training samples and 2264 test samples. Each sample includes recordings of 20 spoken digits transformed into spike sequences using a cochlear model, capturing the rich dynamics of auditory processing. The 700 cochlear model output channels are downsampled to 350 channels.
@@
The DVS gesture dataset Amir et al. (2017) focuses on different hand and arm gestures recorded by a DVS camera. Like the SHD dataset, it has significant temporal structure. It focuses on 11 different hand and arm gestures recorded by a DVS camera. It consists of 1176 training samples and 288 test samples. The $128\times 128$ input frame is downsampled to a $32\times 32$ frame.
CIFAR-10 Krizhevsky (2012) is a widely-used image classification dataset consisting of 60000 32x32 color images across 10 classes, including animals and vehicles. It has no temporal structure. Unlike the other datasets, all input is provided in one single timestep.
@@
Table 6: Network architecture and hyperparameters. The architecture is given as [neurons in hidden layers $\times$ number of hidden layers] - [neurons in output layer].
N-MNIST SHD DVS gesture CIFAR-10
[…]
Timestep size 10 ms 10 ms 20 ms N/A
Batch size 256 32 32 64
Epochs 50 100 70 150
Learning rate 5e-4 7e-4 1e-4 2e-4
Membrane threshold $U_{\text{thr}}$ 0.3 0.3 0.3 0.2
Weight decay constant $\lambda_{W}$ 1e-5 1e-4 1e-5 0
Membrane time constant $\tau_{m}$ 1 ms 100 ms 100 ms N/A
Surrogate steepness $\alpha$ 2 10 10 2
Input spike dropout 0.25 0.2 0.2 0
Forward group size $F$ 8 8 8 512
Refractory dropout 0.8 0.8 0.8 0.7
Momentum noise $\lambda_{\text{MS}}$ 1e-6 0.1 0.1 see A.10
For all four datasets, events of a data point belong to a continuous stream where they have a timestamp and an index position (see figure 8). In the case of N-MNIST, the index corresponds to a position within a $34\times 34$ pixel frame, with each pixel having a binary polarity value (either 1 or 0), leading to a total of $34\times 34\times 2=2312$ distinct input indices. For SHD, the index denotes one of the 350 frequency channels. For DVS gesture, the $128\times 128$ input frame is downsampled to a $32\times 32$ frame. Like N-MNIST, each pixel has a binary polarity value, so in total this gives $32\times 32\times 2=2048$ input indices. Unlike the other datasets, for CIFAR-10, the input events present a continuous value (so they are currents instead of spikes), while also still being assigned an index and timestamp (although the timestamp is irrelevant). For encoding them as inputs to the models the stream of events is sliced in time-bins ($\epsilon_{k},\epsilon_{k+1},$...) along the temporal dimension and the timestamps are discretized to the time-bin index. The time-bins then construct time-frames whereby all the events at each spatial index are accumulated (counted). Each time-frame is then used as input to the model at discrete subsequent timesteps. This is one of the common-place input encoding for SNNs.
Figure 8: Encoding of a continuous temporal stream of input events into frames Hagenaars et al. (2021).
@@ A.8 Results on resource use
A.9 Results on error convergence
Both unlayered training with random scheduling and momentum scheduling achieve lower loss values compared to layered training, as shown in figure 10. At the start of training, random scheduling shows a slightly less steep loss curve, while momentum scheduling shows a steepness comparable to that of layered training. For CIFAR-10, error convergence is shown in figure 12.
Figure 10: Error convergence for different synchronization methods.
A.10 Details on CIFAR-10 dataset with VGG experiments
The details of network structure of the VGG model that we used to train on the CIFAR-10 dataset is shown in figure 11. We have removed (for simplicity) the batch normalization and dropout layers (that explicitly necessitate layer synchronization).
To reduce the memory overhead for training with unlayered backpropagation and make it feasible on an NVIDIA RTX 5000 we provided all the input of each sample in a single timestep (thus eliminating the unfolding across timesteps), and we adopted a forward group size $F=512$ for training and $F=64$ for testing. Additionally we deployed Integrate-and-Fire (IF) neurons that do not leak.
Output predictions were based on the membrane potential accumulated at the output neurons. Because of the very large $F$-group size, MS scheduling policy was difficult to train requiring rather large amounts of annealing noise (the range of values tested are shown in figure 12). Note that a large $F$-group is typically adverse for supporting network asynchrony, but in this experiment we were limited by the training resources given the currently high computational cost of unlayered backpropagation. Nonetheless, the results in figure 7, show that even with such large value range for $F$-group, there is a monotonic trend for accuracy recovery in asynchronous inference, as we decrease $F$.
The training error evolution is reported in figure 12. For momentum scheduling with $\lambda_{\text{MS}}=0.1$ or $\lambda_{\text{MS}}=2.5$, training was unstable after initial convergence and was terminated as soon as the error started to diverge. This was not the case for random scheduling or larger values of $\lambda_{\text{MS}}$.
Figure 11: Simplified VGG network topology without batch normalization and dropout layers.
Figure 12: Training error convergence for CIFAR-10 for different synchronization methods and amounts of momentum noise in the MS scheduling policy.
A.11 Neuromorphic processing and energy efficiency
Neuromorphic chips such as µBrain Stuijt et al. (2021) and Speck Caccavella et al. (2023) are at the forefront of asynchronous spiking neural network (SNN) hardware development, showcasing significant benefits over traditional clock-based architectures. Both chips employ an event-driven, fully asynchronous, architecture, eliminating the need for global or per-layer synchronization. In both cases, all neurons can fire a spike immediately once their membrane potential crosses a specific threshold in response to integrating an incoming current/spike (without waiting for a synchronization signal or an interval timer to expire). Such fully event-driven inference allows independent and instantaneous neuron processing, reducing computational overhead and latency, and directly matches the operational principles of SNNs. However, neurons with fully event-driven asynchronous computation are sensitive to the sub-microsecond timing and exact order dynamics of the incoming spikes, which a time-stepped training algorithm is not sensitive to. As a result, there will be a significant accuracy drop if the deployed SNN of these chips is trained with a time-stepped algorithm.
Traditional synchronous simulation methods for SNNs introduce limitations when mapping to these asynchronous neuromorphic chips. In synchronous simulations, time is discretized into uniform steps, and all neurons in a layer are updated simultaneously, leading to excessive idle computations and artificial delays that are not representative of “real-time” temporal dynamics. In contrast, asynchronous simulation methods align more closely with the intrinsic event-driven nature of neuromorphic hardware. They allow all neurons to react as events occur, mirroring the operational principle of chips like µBrain and Speck. This results in lower latency and more efficient mapping of SNN models onto the hardware.
In the absence of common-place asynchronous training and simulation methods, many other neuromorphic chips (e.g. SpiNNaker Furber et al. (2014), TrueNorth Akopyan et al. (2015), Loihi Davies et al. (2018), Seneca Tang et al. (2023)) resort to explicit layer synchronization primitives (timer-based current integration or explicit signaling) for maintaining consistent performance with synchronous simulated and trained models. Under this constraint these systems execute models event-based, but not end-to-end asynchronously (even though they could). This comes at the cost of increased latency of execution and energy consumption.
Typically, energy consumption on such digital neuromorphic accelerators has two components. A baseline static one that is present even when the system is idle so long as it is powered, and which relates to its hardware attributes and components (memory leakage, clock frequency, manufacturing technology, and other). The second is a dynamic component that relates to the operation of the system for executing a model. The latter has to do with actual computations the system performs, memory I/O (typically the source of the Von Neummann bottleneck), and communication (particularly in multicore systems that entail a network-on-chip). In many architectures, a technique called power-gating helps save substantial energy from static power by switching off parts of the system that are idle.
Model execution is directly related to energy consumption due to dynamic power (e.g. when fetching data from memory and executing arithmetic operations), but also indirectly due to static power (as a function of the inference latency that forces the system to stay powered).